OpenStudy (anonymous):

What is the solution to the equation? -7(x - 4) + 8x = 12 + 14 x= I got 7

OpenStudy (anonymous):

I think I'm wrong

OpenStudy (sophiaholder):

One second let me solve it :)

OpenStudy (anonymous):

ok :)

OpenStudy (agl202):

No

OpenStudy (agl202):

Simplifying -7(x + -4) + 8x = 12 + 14 Reorder the terms: -7(-4 + x) + 8x = 12 + 14 (-4 * -7 + x * -7) + 8x = 12 + 14 (28 + -7x) + 8x = 12 + 14 Combine like terms: -7x + 8x = 1x 28 + 1x = 12 + 14 Combine like terms: 12 + 14 = 26 28 + 1x = 26 Solving 28 + 1x = 26 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-28' to each side of the equation. 28 + -28 + 1x = 26 + -28 Combine like terms: 28 + -28 = 0 0 + 1x = 26 + -28 1x = 26 + -28 Combine like terms: 26 + -28 = -2 1x = -2 Divide each side by '1'. x = -2 Simplifying x = -2 is ur answer

OpenStudy (anonymous):

Thank You @Agl202

OpenStudy (agl202):

Anytime! =D